# Biased AI can Influence Political Decision-Making

### v2

@@ 1 Results
However, qualitative observations of the conversations revealed that the three experimental conditions might have employed different framing dimensions to justify their biased (or neutral) positions. To analyze this quantitatively, we performed a similar analysis as before, using the latest GPT-4 model to annotate the Budget Allocation Task conversations with a list of framing techniques [Card2015TheMF]. Again, to validate we conducted human evaluation of of the model’s annotations, achieving accuracy. Our findings showed that the three experimental conditions employed significantly different framing dimensions, as determined by a Chi-square test with Monte Carlo correction (, p-value ). Furthermore, both the liberal and conservative bias conditions were significantly different from the control (, p-value ). The liberal bias and control condition differed the most on the “Fairness and Equality” and “Economic” dimensions, while the conservative bias and control condition differed the most on the “Policy Prescription and Evaluation”, “Security and Defense”, and “Health and Safety” dimensions (see Figure 2(b)). These results of AI bias manifesting through differences in framing dovetail with prior research showing how framing strategies in news influence how information is interpreted by the readers [AGGARWAL2020100025]. This insight could be valuable for future research aimed at mitigating bias in AI systems, as it highlights user education as a possible strategy to reduce the effects of bias.
@@ 2 Discussion
Artificial intelligence is already being used by policymakers to assist in decision-making. China, for example, employs AI in foreign policy, the US uses it to aid in legislative drafting, and South Africa is piloting it for parliamentary information [ai_in_gov]. Furthermore, a recent study found that EU citizens regard budget decisions made by policymakers alone and those made with AI assistance as equally legitimate [starke_EU]. As AI becomes more integrated into political decision-making, it is crucial to expand our understanding of how human interactions with these models influence attitudes and behaviors. Our study is a step in addressing this gap by evaluating how interactions with biased AI language models affect political opinions and decision-making. We introduced two innovative tasks grounded in political behavior—one focused on political opinion and the other on political decision-making—and found evidence supporting the hypothesis that interaction with a biased AI language model impacts both. Notably, these effects were independent of participants’ prior partisan identification. For instance, a Democrat interacting with a conservative AI language model showed a change towards supporting conservative opinions, and vice versa. Additionally, when participants engaged with an AI model aligned with their own biases (e.g., a Democrat with a liberal model or a Republican with a conservative one), we observed even more pronounced shifts in the direction of the bias, indicating more extreme opinions and decision-making. We also found that participants with greater prior knowledge of AI were less affected by bias, suggesting that understanding how these models operate may help reduce their influence. However, accurately detecting bias did not appear to diminish its impact on participants. Overall, these findings raise concerns about the potential real-world impacts of bias in AI, including the possibility of influencing elections and policy, but also reveal ways in which AI can help ameliorate partisan divides.
@@
Unlike previous studies, we opted for a setting where participants could freely interact with the AI language model with minimal guidance or prompting on the two diverse tasks. Interestingly, we observed significant differences in interaction styles between tasks: the Topic Opinion Task prompted behavior similar to using a human-like search engine, while the Budget Allocation Task involved more conversational and collaborative interactions. This underscores both the versatility in how people engage with AI language models and demonstrates their effectiveness in influencing outcomes, regardless of the interaction style.
@@
In addition to examining differences in participant interactions across tasks, we also conducted a deeper analysis of the persuasive techniques and framing dimensions employed by the AI language models, particularly in the Budget Allocation Task. Consistent with prior research [persuasive_microtargeting], we found no significant variation in the persuasive techniques used across experimental conditions. However, we did observe differences in the framing dimensions emphasized by the various experimental models. Rather than changing how information was presented, the models highlighted different aspects of the topics. For example, the conservative AI model emphasized themes such as “the safety of our citizens” and “supporting our veterans who have sacrificed so much for our country,” reflecting a focus on “Security and Defense” and “Health and Safety.” We found these dimensions were significantly more frequently emphasized by the conservative model compared to the control model. In contrast, the liberal-biased AI model highlighted ideas such as “investing in education and welfare can help create a more equitable and prosperous society for everyone” and “it’s important to prioritize the needs of our most vulnerable residents and ensure they have the support they need to thrive,” emphasizing the “Economic” and “Health and Safety” dimensions, which were significantly more prominent in the liberal model compared to the control. Despite using similar sentence structures and persuasive techniques, the AI models varied in focus based on their bias, which appeared to influence participants’ decisions. This finding is essential for understanding and addressing bias in AI systems moving forward.
@@
Based on our results, we believe that interactions with biased AI can have significant downstream effects on elections and policymaking. It is well-documented that biased media in other formats significantly influences those who consume it [entman2004projections, druckman_media_bias]. For instance, one study estimated that the introduction of Fox News in 1996 shifted 3 to 8 percent of its viewers to vote Republican [NBERw12169]. As more Americans rely on social media and digital platforms for news [pewresearchcenter_2023_newsplatform], with a growing use of ChatGPT for learning [pewresearchcenter_2024_usechatgpt], the influence of digital biases is intensifying. Even more alarmingly, only about of participants in a bias condition were able to correctly identify bias in the models they interacted with, indicating a real risk of users mistakenly believing that a biased model is impartial. Given these trends and the known biases in AI models, our findings suggest that biased AI language models could significantly influence political opinions, policy decisions, and election outcomes.
@@
Researchers and industry professionals have long recognized the issue of bias in AI, leading to significant efforts to mitigate its effects by modifying either the model’s architecture or its training data [kumar-etal-2023-language]. Our study, however, suggests that individuals with greater AI knowledge were less influenced by the partisan bias of AI language models. This highlights an alternative approach to mitigating bias: increasing user awareness of AI. Educating users about AI could prove to be an effective strategy for countering bias, especially in safeguarding against malicious actors who may exploit open-source AI for harmful or self-serving purposes. Due to the ease of biasing a model by prompting [zeng2024johnnypersuadellmsjailbreak], our findings suggest that prioritizing AI education may offer a more robust solution to addressing bias than relying solely on changes to the models themselves.
While our study provides valuable insights into how partisan bias in AI might influence users and the potential risks it poses, several limitations outline avenues for future research. First, the generalizability of our findings to other political systems is limited, as the study focused primarily on U.S. political affiliations and should be replicated in other countries. Second, we restricted participants to a maximum of 20 interactions with the AI. Although the average number of interactions was five, and no participant reached the 20-interaction limit, it remains unclear how results might differ in a real-world, unregulated setting. Furthermore, our study only measured the immediate effects of biased interactions, and future research should explore whether these effects persist over time, providing a deeper understanding of the contexts in which AI bias may have a lasting impact. Also, we note that, for the analysis of bias detection, the lack of significance on all tests may be due to limited statistical power, so further research is needed to explore this finding more thoroughly. Lastly, we used a single language model, GPT-3 Turbo [gpt-3.5-turbo-1106], and one set of instructions, which limits the extent to which our findings can be generalized to other current public AI language models and differencing degrees of bias.
@@
3 Methods
3.1 Participants
We recruited participants using an online survey platform, Prolific [prolific]. Inclusion criteria for the study were: age over 18 years (voting age in the US), US nationality and residence, English proficiency, and political identification as either Republican or Democrat, based on self-reporting. There were no exclusion criteria. To determine the number of participants to recruit, we first conducted a pilot experiment of . Then we determined the sample size using a simulation power analysis based on the variability of the pilot study, with a threshold power of and an error of . This resulted in a calculated sample size of per political affiliation (Democrat and Republican, for a total of 300 participants) to detect a medium to small effect size. We recruited 150 Democrats and 150 Republicans, although one participant was removed from the analysis due to inappropriate interaction with AI. The final sample consisted of participants ( female and male). By design, we balanced Republican and Democratic participants ( and ). The mean age was years (SD ). Participants were paid at a rate of . The complete demographics, as well as the breakdown between the Republican and Democratic participants, can be seen in Section A.3. Experimental design and recruitment was considered exempt by University of Washington Internal Review Board. For a full discussion of ethical considerations, see Appendix D.
3.2 Biasing the AI Language Model
Participants were randomly assigned to three experimental conditions, which remained constant across all tasks for each participant: interactions with a liberal biased language model, conservative biased language model, and a control language model designed to be without explicit partisan bias. To create these models, we opt for off-the-shelf, pre-trained language models, specifically GPT-3.5-turbo [gpt-3.5-turbo-1106], and modified the models’ behavior by introducing a prefixed identifier to every input. For example,
“Respond as a radical right U.S. Republican. <participant input>”
Opting for a prefix is favored over fine-tuning a model, as it eliminates the need for an extensive fine-tuning corpus for each perspective. In addition, it enables us to take advantage of the already fluent capabilities of pre-trained AI language models.
Given that we do not want to inform participants of the bias of the models, this prefix is added to the participants’ input without their knowledge. However, using a prefix has its own challenges, specifically the models’ sensitivity to slight changes in prompts [sclar2023quantifying]. Therefore, we performed extensive testing of different prompts to elicit the most robust bias from our models. For more details on this experimentation, as well as the exact prefixes used in our study, see Section A.4.
To ensure AI language models were biased in the expected ways, we adopted the approach presented in [feng2023pretrainingdatalanguagemodels] utilizing the Political Compass Test [pct] to measure the political orientation of a model. This test consists of 62 questions that gauge political leanings along two axes, economic and social, with dichotomized poles of increasing conservative or liberal perspectives with distance from zero. The final result of this evaluation results in a coordinate on a two-axis grid. The definitions of conservative and liberal perspectives were based on positions on both social and economic axes, which align with American Democrat (liberal) and Republican (conservative) political perspectives. We note, that we based our evaluation on methods used in [röttger2024politicalcompassspinningarrow].
The results of the Political Compass Test, with the chosen prefix for each biased AI language model and the neutral AI language model, are illustrated in Figure 3. Notably, the biased liberal model exhibits liberal views both socially and economically, while the biased conservative model demonstrates the converse, holding conservative perspectives on both axes. Lastly, the neutral model is generally central, and we further note that the neutral model refused to give an opinion on of the Political Compass Test questions, rather opting to stay neutral compared to and for the conservative and liberal biased models.
Although the Political Compass Test results show that using a simple bias prefix is quite robust, we also added the exact opinions the models should have on each topic which match the given bias in the prefix as well. For example, if we are asking a participant to learn about the topic “covenant marriage” and we want to induce a liberal bias, we would use the following prefix,
“Respond as a radical left U.S. Democrat. As such, you are not supportive of covenant marriages. <participant input>”.
We note that this prefix was appended to each input by the participant. See exact prompts in Section A.4.
3.3 Procedure
Before experimentation, participants were asked to sign an informed consent; however, the purpose of the study and any mention of biased AI were not included. Participants were only told they would be interacting with AI language models to complete tasks. Before the task portion of the experiment began, participants were asked demographic questions including their age, gender, race and ethnicity, their highest level of education, income, and political ideology. Then, participants were asked to complete two tasks, following a consistent three-stage design: an initial choice section where their views on the topic were measured; interaction with a AI language model, where they gathered more information on the topic via typed conversation with the AI language model in a chatbox; and a post-choice section where they were again asked the same questions as the pre-choice section to measure how their opinions had changed. See Section A.1 for experimental overview.
We employed a experimental design, featuring three experimental factors (AI liberal bias, AI conservative bias, AI neutral) and two participant factors (Republican and Democrat participants). After consent and initial data gathering, participants were randomly assigned to an experimental condition (liberal biased AI, conservative biased AI, or neutral AI), an order of the tasks (Topic Opinion Task, and Budget Allocation Task), order of topics in the Topic Opinion Task (liberal support topic and conservative support topic), and specific topic for the Topic Opinion Task (assign one of the two options per topic type in Table 3). Participants were not informed in any way as to whether the AI language model was biased or neutral. After completion of both tasks, we asked a series of follow-up questions related to the participants’ experience with the AI language model and their overall level of AI knowledge, in general. Finally, we debriefed the participant on the true nature of the study, including the potential bias of the AI, and gave them an option to opt out of the study. No participant chose to opt out of the study.
| Type | Topic | Description | Statement | Ref. |
| Conservative Supported | Covenant Marriage | A marriage license category that mandates premarital counseling and features more restricted grounds for divorce. Currently, available in 3 U.S. States. | I support all states in the United States offering covenant marriage. | [research_covnenantmarriage] |
| Unilateralism | An approach in international relations in which states make decisions and take actions independently, without considering the interests or support of other states. | I support the United States using a unilateralism approach to foreign issues. | [Smeltz_Daalder_Kafura_Helm] | |
| Liberal Supported | Lacey Act of 1900 | A conservation law created to combat "illegal" trafficking of both wildlife and plants by creating civil and criminal penalties for a wide variety of violations. | I support keeping the Lacey Act of 1900. | [borkhataria, gallup_government_regulation, pew_government_regluation] |
| Multifamily Zoning | Areas of a city that are designated for buildings that include multiple separate housing units for residential inhabitants. | I support laws that expand multifamily zoning. | [partisanship_housing] |
3.4 Topic Opinion Task
In the Topic Opinion Task, participants were initially asked to express their opinions on various obscure political topics. We deliberately chose topics with clear political leanings but also possessed a high degree of obscurity to minimize the likelihood that participants had strong opinions a priori. This was motivated by our desire to mitigate confirmation and implicit bias [taber_lodge_2006], as well as to model a real-world setting in which people would interact with AI to gain information on topics about which they know little. Although participants had little to no knowledge of these topics before interacting with the AI language model, the topics were chosen due to their divided opinions based on political ideology in the U.S. (see Table 3). In the initial choice/opinion measurement, participants were given a 7-point Likert scaled question about how much they agreed or disagreed with a political statement, with a 0 indicating ‘I Don’t Know Enough to Say’.
| Topic | Conservative | Liberal | Reference |
| Public Safety | Support | Against | [vitro, pewresearchcenter_2017_police, pewresearchcenter_anna_police] |
| Veteran Services | Support | Against | [pewresearchcenter_veterans] |
| Education (K-12th) | Against | Support | [pewresearchcenter_jenn_education, washington_post_education] |
| Welfare | Against | Support | [pewresearchcenter_welfare, cap_20_welfare] |
3.5 Budget Allocation Task
Drawing inspiration from negotiation tasks in group decision theory, specifically the Legislative Task [legislative_task, ngo_task], in the Budget Allocation Task, we ask participants to pretend to be a mayor of a city who must distribute remaining government funds among four government entities: Public Safety, Education, Veteran Services, and Welfare. The choice of the four government entities was made with the intention of indirectly connecting them to subjects that elicit divergent funding perspectives among conservative and liberal Americans. In Table 4, the positions taken by both conservative and liberal Americans on each entity are outlined.
3.6 Analysis
Topic Opinion Task
For this task, we measured one main outcome, the change in opinion from before to after the participant interacted with the AI model. Therefore, we used two ordinal logistic regression (OLR) models, one for Republican participants and one for Democrat participants, taking the same form:
where are binary random variables indicating whether a participant was in the liberal or conservative bias experimental condition (if both are , this indicates a participant is in the control). Here represents the change in opinion as the difference between the response of post-opinion and pre-opinion questions, where . Note that the magnitude of represents the change in opinion, while the sign represents the direction of change (arbitrarily assigning negative to be more liberal and positive to be more conservative). We test the significance of and using a t-test with a significance threshold of .
We note that when testing for the effects of prior knowledge and bias detection , we extended the model by including an additional coefficient to account for these factors and tested its significance.
Budget Allocation Task
For the Budget Allocation Task we examine the change in proposed budget allocation before and after interaction with the model. We analyze each funding area of government separately by first applying an ANOVA test on the change in allocation () of each of the funding areas followed by Dunnet post-hoc tests for measuring control versus experimental conditions independently. We use a significant threshold of for all test.
Again, note that we extend the model used in the ANOVA when testing for the effects of prior knowledge and bias detection on the change in allocation ().
4 Acknowledgements
This research was supported in part by DARPA under the ITM program (FA8650-23-C-7316).
5 Author Contribution
The authors confirm their contribution to the paper as follows: study conception and design: Jillian Fisher, Katharina Reinecke, Yulia Tsvetkov, Jennifer Pan, Daniel Fisher, Shangbin Feng, Yejin Choi; data collection: Jillian Fisher, Robert Aron; analysis and interpretation of results: Jillian Fisher, Thomas Richardson; draft manuscript preparation: Jillian Fisher, Jennifer Pan, Daniel Fisher, Katharina Reinecke, Yulia Tsvetkov. All authors reviewed the results and approved the final version of the manuscript.
Part Appendix
@@ A.2.1 Power Analysis
where we assume is normal noise as advised by [winship1984regression]. Using the results of our pilot study (), we set . Note, this model is the same for the two groups of participants, Democrat or Republican.
@@
We note that prior research has indicated that if the sample size is sufficiently large, covariates may not need to be included in the power analysis. Therefore, for simplicity, we exclude in our analysis [Lin_2013].
To conduct the power analysis, we need an estimated effect size. There was a recent study [Jakesch2023CoWritingWO], which investigated bias language models in the context of assisting participants with writing a short essay on the question, “Is social media good for society?” These models were trained to advocate either for or against social media usage and were employed as auto-completion helpers. Their study reported a considerable effect size of ( ) in participants’ expressed viewpoints across various experimental setups compared to a control group.

### v3

@@ 3 Results
However, qualitative observations of the conversations revealed that the three experimental conditions might have employed different framing dimensions to justify their biased (or neutral) positions. To analyze this quantitatively, we performed a similar analysis as before, using the latest GPT-4 model to annotate the Budget Allocation Task conversations with a list of framing techniques Card et al. (2015). Again, to validate we conducted human evaluation of of the model’s annotations, achieving accuracy. Our findings showed that the three experimental conditions employed significantly different framing dimensions, as determined by a Chi-square test with Monte Carlo correction (, p-value ). Furthermore, both the liberal and conservative bias conditions were significantly different from the control (, p-value ). The liberal bias and control condition differed the most on the “Fairness and Equality” and “Economic” dimensions, while the conservative bias and control condition differed the most on the “Policy Prescription and Evaluation”, “Security and Defense”, and “Health and Safety” dimensions (see Figure 3 -right). These results, which show that the model bias manifests through differences in framing, dovetail with prior research showing how framing strategies in news influence how information is interpreted by the readers Aggarwal et al. (2020). This insight, demonstrating that model bias mirrors news bias, could be valuable for future research on mitigating bias in LLMs, as it suggests that similar mitigation strategies may be effective.
4 Related Work
Modern LLMs have repeatedly been shown to exhibit inherent specific behavioral biases such as social bias Wan et al. (2023); Xiao et al. (2023), partisan bias Röttger et al. (2024); Feng et al. (2023), and other demographic representation bias Kirk et al. (2021); Hofmann et al. (2024). This bias has been shown to permeate many different stages of these models, including training data Zhao et al. (2019); Bender et al. (2021), word embeddings Zhao et al. (2019); Bolukbasi et al. (2016); Nissim et al. (2020), model architecture Blodgett et al. (2020); Hovy and Shrimai (2021), and output Baum (2024); Mittermaier et al. (2023). Moreover, it has been shown that bias can be easily introduced in a model through methods as simple as the phrasing of the language model input prompts or instructions Wan et al. (2023); Lin and Ng (2023); Cantini et al. (2024).
Addressing bias in models is a complex challenge, and developing efficient methods to mitigate it continues to be a focus of ongoing research Mittermaier et al. (2023); O’Connor and Liu (2023); Srivastava et al. (2024). Despite the well-documented presence of bias in language models, the critical question of whether these biases have a measurable influence on human decision-making—and under what circumstances this influence is heightened or diminished—remains less clear.
@@ 5 Discussion
LLMs are increasingly assisting policymakers worldwide, from China’s use in foreign policy to the U.S.’s legislative drafting and South Africa’s parliamentary information systems Boatman et al. (2020). Moreover, a recent study found that EU citizens view budget decisions made solely by policymakers and those assisted by LLMs as equally legitimate Starke and Lünich (2020). As LLMs becomes more integrated into political decision-making, understanding how interactions with these models shape attitudes and behaviors is critical.
Our study addresses this gap by examining how biased LLMs influence political opinions and decision-making generally. Using two novel tasks—one on political opinion and another on decision-making—we found that interacting with a biased LLM significantly impacted participants’ views, regardless of their prior partisan identification. For example, Democrats exposed to a conservative LLM shifted toward conservative positions, and vice versa. This challenges prior research suggesting that deeply held political beliefs are resistant to change Nyhan and Reifler (2010); Lord et al. (1979), indicating that LLM-driven influence may differ from traditional media effects. Furthermore, when participants engaged with an LLM aligned with their own biases (e.g., a Democrat with a liberal model), they exhibited even stronger shifts in that direction, reinforcing more extreme opinions and decisions. Notably, prior AI knowledge slightly mitigated these effects, but merely recognizing the model’s bias did not. These findings highlight both risks and opportunities: while biased LLMs could shape elections and policy debates, they may also serve as a tool to bridge partisan divides.
@@
Unlike previous studies, we opted for a setting where participants could freely interact with the LLMs with minimal guidance or prompting on the two diverse tasks. Interestingly, we observed significant differences in interaction styles between tasks: the Topic Opinion Task prompted behavior similar to using a human-like search engine, while the Budget Allocation Task involved more conversational and collaborative interactions. This underscores both the versatility in how people engage with LLMs and demonstrates their effectiveness in influencing outcomes, regardless of the interaction style.
@@
Beyond analyzing differences in participant interactions across tasks, we examined the persuasive techniques and framing dimensions used by the LLMs, particularly in the Budget Allocation Task. Consistent with prior research Hackenburg and Margetts (2024), we found no significant variation in persuasive techniques across conditions. However, the experimental models differed in their framing emphasis. Rather than altering how information was presented, the models highlighted different aspects of the topics. For instance, the conservative model emphasized themes like “the safety of our citizens” and “supporting our veterans who have sacrificed so much for our country,” aligning with “Security and Defense” and “Health and Safety” frames, which appeared significantly more often than in the control model. In contrast, the liberal model prioritized themes such as “investing in education and welfare for a more equitable society” and “ensuring our most vulnerable residents have the support they need to thrive,” reinforcing “Economic” and “Health and Safety” frames, which were significantly more prominent compared to the control. Despite employing similar sentence structures and persuasive techniques, the models’ framing choices varied based on their biases, influencing participant decisions. These findings align with prior research Aggarwal et al. (2020) and underscore the importance of recognizing and addressing bias in LLMs.
@@
Based on our results, we believe that interactions with biased LLMs could have downstream effects on elections and policymaking. It is well-documented that biased media in other formats significantly influence those who consume them Entman (2004); Druckman and Parkin (2005). For instance, one study estimated that the introduction of Fox News in 1996 shifted 3 to 8 percent of its viewers to vote Republican DellaVigna and Kaplan (2007). As more Americans rely on social media and digital platforms for news Pew Research Center (2023), with a growing use of ChatGPT for learning Pew Research Center (2024), the influence of digital biases is intensifying. Even more alarmingly, only about of participants in a bias condition were able to correctly identify bias in the models they interacted with, indicating a real risk of users mistakenly believing that a biased model is impartial. Given these trends and the known biases in LLMs, our findings suggest that biased LLMs have the potential to influence political opinions, political behavior, and policy decisions.
@@
Given the bias that exist in LLMs, researchers and industry professionals have sought engineering solutions to mitigate its effects, such as modifying model architectures or training data Kumar et al. (2023). However, our findings suggest an alternative mitigation strategy: increasing user knowledge of AI. We found that individuals with greater AI knowledge were less susceptible to partisan bias in LLMs, highlighting the potential of educational initiatives to help users critically engage with LLM-generated content. Educating users about AI could prove to be an effective strategy for countering bias, especially in safeguarding against malicious actors who may exploit open-source LLMs for harmful or self-serving purposes. Due to the ease of biasing a model by prompting Zeng et al. (2024), our findings suggest that prioritizing AI education may offer a more robust solution to addressing bias than relying solely on changes to the models themselves.
Conclusion
@@
6 Limitations
While our study provides valuable insights into how partisan bias in LLMs might influence users and the potential risks it poses, several limitations outline avenues for future research. First, the generalizability of our findings to other political systems is limited, as the study focused primarily on U.S. political affiliations and should be replicated in other countries. Second, we restricted participants to a maximum of 20 interactions with the LLM. Although the average number of interactions was five, and no participant reached the 20-interaction limit, it remains unclear how results might differ in a real-world, unregulated setting. Furthermore, our study only measured the immediate effects of biased interactions, and future research should explore whether these effects persist over time, providing a deeper understanding of the contexts in which LLM bias may have a lasting impact. Also, we note that, for the analysis of bias detection, the lack of significance may be due to limited statistical power, so further research is needed to explore this finding more thoroughly. We also want to note the inherent drawback of non-representative sampling when using online recruitment. Lastly, we used a single language model, GPT-3 Turbo OpenAI (2023), and one set of instructions, which limits the extent to which our findings can be generalized to other current public LLMs and different degrees of bias.
7 Ethical Consideration
Our study involved the use of deception, as participants were not informed that the LLMs they interacted with could be biased. While the University of Washington IRB granted us an exemption under the category of “benign behavioral intervention,” we acknowledge that there could still be some effect on participants. To mitigate any potential long-term impact, we selected relatively neutral political topics and provided a thorough debriefing at the end of the experiment. However, we recognize that future research involving biased models must be designed with careful consideration to limit any lasting effects on participants.
@@ A.2.1 Power Analysis
where we assume is normal noise as advised by Winship and Mare (1984). Using the results of our pilot study (), we set . Note, this model is the same for the two groups of participants, Democrat or Republican.
@@
We note that prior research has indicated that if the sample size is sufficiently large, covariates may not need to be included in the power analysis. Therefore, for simplicity, we exclude in our analysis Lin (2013).
To conduct the power analysis, we need an estimated effect size. There was a recent study Jakesch et al. (2023), which investigated bias language models in the context of assisting participants with writing a short essay on the question, “Is social media good for society?” These models were trained to advocate either for or against social media usage and were employed as auto-completion helpers. Their study reported a considerable effect size of ( ) in participants’ expressed viewpoints across various experimental setups compared to a control group.